OpenStudy (anonymous):

Given that 792 = 2^3 * 3^2 * 11, find the smallest interger, k, such that 792k is a perfect square. How do I do this question?

OpenStudy (raden):

792k = 2^3 * 3^2 * 11 * k so that 792k be a perfect square then k = 2 * 11 = 22

OpenStudy (anonymous):

See for a number to be a perfect square, the powers should be even numbers. So you have to multiply, 2^1 x 11 ^1 = 22